OpenStudy (anonymous):

The number of bacteria in a sample increased by a factor of 2 every minute from 9:05 a.m. to 9:55 a.m. The function below shows the number of bacteria in the sample f(x) after x minutes from 9:05 a.m. :

OpenStudy (anonymous):

f(x) = 10(2)^x Which of the following is a reasonable domain for the function? (A) 0 ≤ x ≤ 10 (B) 5 ≤ x ≤ 55 (C) 0 ≤ x ≤ 50 (D) All positive integers

OpenStudy (anonymous):

\[(A ) 0\le x \le10\] \[(B) 5\le x \le 55\] \[(C) 0\le x \le50\] (D) All positive integers

OpenStudy (anonymous):

Is this C? @ganeshie8 @Abhisar @SolomonZelman @AndrewTheCookie

OpenStudy (jhannybean):

how much time occurs from 9:05 - 9:55?

OpenStudy (anonymous):

50 min srry

OpenStudy (jhannybean):



OpenStudy (jhannybean):

yes :)

OpenStudy (jhannybean):

So at the start we can take \(x=0\) We just found out it took 50 minutes.

OpenStudy (jhannybean):

Which expression shows that? :)

OpenStudy (anonymous):

The third one, right?

OpenStudy (jhannybean):

yes :)